# Skill vs. Chance Quantification for Popular Card & Board Games

### v1

@@ 4.2 Rapid, one year data
In the version of Chess, there is a broader format that allows a larger number of players to play against each other. Now we first use only one year, a single edition of the tournament data first. Since no. of games is $<$ 13 for some players (maybe they quit after some rounds), took both half game count as predictor. Dropped second half as it’s not significant. See the result in table 3. Now the explanatory factors return significant values with ratings having a positive effect but performance in the first half becoming negative. This could be due to the correlation between performance and rating. To understand this better, we move to a two year analysis next.
@@ 4.3 Rapid, two year data
In this extended version of the analysis using two years’ data, we have a total of 96 players common in both editions of the tournament. Initial and final ratings denote ratings in 2022 and 2023 respectively. Response is $\phi^{-1}$(win proportion in 23). As one can see from the result in table 4, now performance, initial rating and second year rating all has the correct sign and are significant.
@@
4.4 Chess Online Data
For an empirical investigation of the robustness of our findings, we need a tournament setting where a large pool of players play against each other. These results can be used to empirically define the variability (or otherwise) of our results using resampling techniques.
4.4.1 Resampling Procedure
We have applied the idea of empirical bootstrap to the regression problem. In this case, the empirical bootstrap is also called paired bootstrap. Given the original sample $\left(X_{1},Y_{1}\right),\cdots,\left(X_{n},Y_{n}\right)$, we generate a new sets of IID observations
$\left(X_{1}^{*},Y_{1}^{*}\right),\cdots,\left(X_{n}^{*},Y_{n}^{*}\right)$
such that for each $\ell$,
$P\left(X_{\ell}^{*}=X_{i},Y_{\ell}^{*}=Y_{i}\right)=\frac{1}{n},\quad\forall i=1,\cdots,n$
Namely, we treat $\left(X_{i},Y_{i}\right)$ as one object and we sample with replacement $n$ times from these $n$ objects to form a new bootstrap sample. Thus, each time we generate a set of $n$ new observations from the original dataset. Assume we repeat the entire process $B$ times, we would obtain
$\begin{gathered}\left(X_{1}^{*(1)},Y_{1}^{*(1)}\right),\cdots,\left(X_{n}^{*(1)},Y_{n}^{*(1)}\right)\\ \left(X_{1}^{*(2)},Y_{1}^{*(2)}\right),\cdots,\left(X_{n}^{*(2)},Y_{n}^{*(2)}\right)\\ \vdots\\ \left(X_{1}^{*(B)},Y_{1}^{*(B)}\right),\cdots,\left(X_{n}^{*(B)},Y_{n}^{*(B)}\right)\end{gathered}$
$B$ sets of fitted coefficients. We then estimate the variance by
$\widehat{\operatorname{Var}}_{B}\left(\widehat{\beta}_{0}\right)$ $=\frac{1}{B}\sum_{\ell=1}^{B}\left(\widehat{\beta}_{0}^{*(\ell)}-\bar{\beta}_{0}^{*}\right)^{2},\quad\bar{\beta}_{0}^{*}=\frac{1}{B}\sum_{\ell=1}^{B}\widehat{\beta}_{0}^{*(\ell)}$
$\widehat{\operatorname{Var}}_{B}\left(\widehat{\beta}_{1}\right)$ $=\frac{1}{B}\sum_{\ell=1}^{B}\left(\widehat{\beta}_{1}^{*(\ell)}-\bar{\beta}_{1}^{*}\right)^{2},\quad\bar{\beta}_{1}^{*}=\frac{1}{B}\sum_{\ell=1}^{B}\widehat{\beta}_{1}^{*(\ell)}$
[…]
@@ 4.4.2 Regression Results
For this study, data has been collected from Lichess, an online chess platform. Data of blitz games of March and April 2024 have been used, and only games played by players rated 2300+ on Lichess have been considered. Further, only the players having played at least 5 games in each month have been considered for the study. The database consists of data for 5267 such players.
Since Lichess rating is updated after every game(unlike FIDE tournaments, where rating is not updated between tournament games), the experience gained in the months of March and April (considered to be the two halves) have been quantified in the following way-
@@
  • Second Half Experience = $\frac{m_{2}+m_{3}}{2}-m_{1}$, where $m_{2}$ and $m_{3}$ are the ratings after the first and last matches in April respectively. (Note: players might have unrecorded matches between the calculation of $m_{1}$ and $m_{2}$, since the data only contains games where both players are rated 2300+ .)
It is to be noted that Lichess ratings and FIDE ratings are calculated slightly differently, and we perform origin-scale transformation on the experiences before using them for regression. Both the quantities used to quantify the experience in either half have been scaled to a range of 0 to 10.
The regression summary is provided in the following table-
@@ 4.4.3 Bootstrap Results
For performing resampling, 10000 bootstrap samples were drawn with replacement from the Lichess online database used for the study, each having size same as the original dataset. The regression was performed on all of the resampled datasets, and the bootstrap mean, variances and $95\%$ confidence intervals for the regression coefficients are provided in the following table.
@@
5 Two player Rummy
For the analysis of this card game we use actual data from online games. We use data for 20,000 players, results are in table 7. The results show strong impact of experience in the face of performance. Current experience also has some impact (t-statistic being more than 1).
@@ 5 Two player Rummy
Residual standard error: 0.2279, F-statistic: 88.33 on (3, 19996) df, p-value: 0.0000
5.1 Resampling results
@@
6 Analysis for Ludo
@@ 6.1 Ludo Experimental Data
For Ludo, we have used pre-launch experimental data with a small number of players (see table 9. The regression results are presented in table 10. Here previous performance plays an important role, in addition we also see the number of games in the second tranche being a serious factor with a negative sign. This is probably due to fatigue, which we will investigate in more detail elsewhere. The positive contribution of previous experience is also mildly important.
Player ID 1st Half Win % 2nd Half Win % $\phi^{-1}$(1st Half Win %) $\phi^{-1}$(2nd Half Win %) 1st Half Games 2nd Half Games
@@ 6.2 Ludo Online Data
This study uses data obtained from a new variant of the Ludo game as an integrated offering on the My11Circle (“MEC”) platform, a fantasy gaming platform. The primary objective of launching the Ludo game on the MEC platform has been to collect the real user game play data. We use this data to conduct our Study. The dataset consists of a total of 265062 players, who have played a total of 3033645 games among themselves during $5^{th}$ June to $23^{rd}$ July. The subsetted dataset used for the actual analysis includes players who have played at least 10 games in both months, consists of a total of 16197 players, who have played a total of 2018935 games, and 1191656 games among themselves during $5^{th}$ June to $23^{rd}$ July.
@@
The regression results firmly show that the role of previous experience (learning) and previous performance (innate skill) is quite prominent in determining current performance of a player. This is a strong indicator of the importance of skill in this Ludo game.
6.2.1 Bootstrap Results
To understand the robustness of our results, we now perform a 10000 fold bootstrap analysis of the regressions results.
@@ 6.2.1 Bootstrap Results
The consistency of the results through the bootstrap analysis further strengthen our conclusions.
7 Teen Patti
We analyse the two most popular variations of this game, No Limit (result is in table 13) and Regular (result is in table 14). We also consider 5 player games only, as these are the most common. In both the versions, previous performance turns out to be an important factor which probably is also related to innate skills. The learning feature is supported by data as previous experience has a negative impact on current performance.
@@ 7 Teen Patti
7.1 Resampling Analysis for Teen Patti
7.1.1 No Limit
@@
Table 15: Bootstrap regression summary for 10000 bootstrap estimates
7.1.2 Regular
@@
Table 16: Bootstrap regression summary for 10000 bootstrap estimates
8 Overall Discussion
@@ 8.1 Qualitative Considerations
In this report, our main aim is to draw comparative statements regarding the effect of Luck vs. Skill in the popular card and board games considered (tables 4 to 14). The initial discussion is around the importance of ”innate skill” (performance in general) and ”learning” (experience: learning by doing). Looking at the comparative results in table 17 we can assess this for the games under consideration. We certainly look at Chess as the benchmark case. Here rating proxies for skill acquired through experience. The regression results are very clear on the role of previous performance and rating.
Among the other three games, first consider Teen Patti, as this is popularly considered to be the “least skilled” game. Here we see a significant role of “innate skill” but no learning effect. This skill may be attitudinal (bluffing) or cognitive (recognising the behavioural markers of the opponent). But, there is no strategic learning.
Moving on to Rummy the results are very different. Again the relevant ”innate skill” here is the choice of strategy in selection and retention of cards, and there is a significant learning effect. In fact this is true even for current experience (albeit at a lower level of significance). Contrasted with this, if we consider Ludo, we again see both skill as well as learning effect but the effect is now slightly weaker. In fact, the experience effect in Ludo is somewhat mixed with a positive and negative sign.
@@
We see that performance in all the games under consideration rely on innate skill (player heterogeneity), which is one indicator of skill base. Interim findings also suggest that an one shot game like Teen Patti is less amenable to learning compared to the repeated interaction games, e.g. Rummy and the two board games considered. Among the board games, Ludo illustrates a slightly weaker effect of learning compared to Rummy and, of course, Chess is the most clear cut skill based game. Based on our limited analysis, we can tentatively rank the four games in order of increasing skill and learning dependence as
@@ 8.2 The Skill Score
We now attempt a scoring exercise on a $[0,1]$ scale for “Skill”. This is tentative as it requires several subjective choices and can be confirmed only when we have more detailed game data on Ludo and Chess.
@@
$x_{i}=\left\{\begin{array}[]{lll}0&if&t_{i}<0\\ t_{i}/4&if&0\leq t_{i}\leq 2\\ (t_{i}+1)/6&if&2<t_{i}\leq 5\\ 1&if&5<t_{i}\end{array},i=1,2,3.\right.$
@@
Finally to define a skill score we propose that past experience should have twice the importance as innate skill, to give learning a higher weightage. In turn, innate skill in the form of past performance should also receive twice as much weight as current experience as current experience may be confounding learning and fatigue. Thus, we define the skill score as
$s=\frac{2x_{1}+4x_{2}+x_{3}}{7}\in[0,1].$
Of course, this formulation is just one possible among many and, end of the day, is a matter of subjective choice. We hope that we can objectify this choice with a more comprehensive analysis in future work.
Below we illustrate with the games considered here in table 18. The Scores obtained almost confirms our ranking posited above, the only exception being Ludo scores slightly higher than Rummy now. As these two numbers are very close to each other, we think that the results may be sensitive to the choice of cut-offs and weightages in the score formula.
@@ 8.3 Resampling Results on Skill Score
The results are not only dependent on the choice of the score function. Sampling fluctuation inherent in the data may also influence our scores and hence ranking. So we also carry out a bootstrap analysis with our regression model. We will get a range of t-stat based on the range of regression coefficients from bootstrap analysis. This will give us a range of $t_{i}$ values and hence a range of $x_{i}$ values. Thus, instead of a point estimate for the score function $s$ for each game, it will have a range of values. We can then compare distribution of the score values to arrive at a comprehensive ranking template. Details follow.
@@
For each of the games we use 10000 bootstrap samples to generate the t-statistics. Hence, we get a distribution of scores for each of the combinations in table 19. These are presented in figures 6 for Chess, 7 for Rummy, 8 for Ludo and finally 9 for Teen Patti.
@@
Figure 6: Distribution of Scores for Various Transformations for Chess Bootstrapped Data
Figure 7: Distribution of Scores for Various Transformations for 2-Player Rummy Bootstrapped Data
Figure 8: Distribution of Scores for Various Transformations for Ludo Bootstrapped Data
Figure 9: Distribution of Scores for Various Transformations for Teen Patti Bootstrapped Data
To simplify our discussion, we produce the median and an 80% confidence interval for all the score distributions in table 21 where the degeneracy of Chess score is clear. It also clear that Teen Patti is less skilled than either Rummy or Ludo (the confidence intervals are practically separated) who are in turn less skilled than Chess. The comparison between Rummy and Ludo is unclear, as already guessed at earlier. The distributions overlap with Ludo having a little more spread and the centre of the distributions (median) also switching order between Rummy and Ludo for alternative parametric configurations. This clearly indicates that these two games are very similar in total skill content as indicated earlier. Of course, the kind of skill required may be quite diverse.
@@
8.4 Concluding Remarks
@@
The conclusions are quite nuanced with Chess being the most skilled (extreme) and Teen Patti on the other end (but not an extreme score, around 0.45). While Rummy and Ludo are in between, the ranking between them is ambiguous. Depending on cutoffs and weightage schemes, the score distribution is always overlapping for this two with a wider range for Ludo than Rummy (possibly due to the larger sample of Rummy data).

### v2

@@ 4.2 Rapid, one year data
In the version of Chess, there is a broader format that allows a larger number of players to play against each other. Now we first use only one year, a single edition of the tournament data. Since the number of games is $<$ 13 for some players (maybe they quit after some rounds), we consider both half game counts as predictors and then dropped the second half as it’s not significant (see the result in table 3). Now the explanatory factors return significant values (p values being less than 0.05) with ratings having a positive effect, but performance in the first half becomes negative. This could be due to the correlation between performance and rating. To understand this better, we move to a two-year analysis next.
@@ 4.3 Rapid, two year data
In this extended version of the analysis using two years’ data, we have a total of 96 players common in both editions of the tournament. Initial and final ratings denote the ratings in 2022 and 2023 respectively, and the response is $\phi^{-1}$ (win proportion in 23). The results in Table 4 show that performance, initial rating, and second-year rating all have positive coefficients, which aligns with the expectation that both prior ratings and past performance significantly influence current outcomes (with p-values less than 0.05)
@@
4.4 Online Chess Data
@@ 4.4.1 Regression Results
For this study, data has been collected from Lichess (https://lichess.org/), an online chess platform. Data of blitz games from March and April 2024 have been used, and only games played by players rated 2300+ on Lichess have been analyzed due to the availability of publicly accessible detailed data as opposed to players with lower ratings from whom public data is not available. Further, only the players who have played at least 5 games in each month have been considered for the study. The database consists of data for 5267 such players.
Since Lichess rating is updated after every game (unlike FIDE tournaments, where rating is not updated between tournament games), the experience gained in the months of March and April (considered to be the two halves) has been quantified in the following way.
@@
  • Second Half Experience = $\frac{m_{2}+m_{3}}{2}-m_{1}$, where $m_{2}$ and $m_{3}$ are the ratings after the first and last matches in April respectively.
  • Note that the players might have unrecorded matches between the calculation of $m_{1}$ and $m_{2}$, since the data only contains games where both players are rated 2300+.
It is to be noted that Lichess ratings and FIDE ratings are calculated slightly differently, and we perform origin-scale transformation on the experiences before using them for regression. Both the quantities used to quantify the experience in either half have been scaled to a range of 0 to 10. The regression summary is provided in table 5.
@@ 4.4.2 Bootstrap Results
For performing resampling, 10000 bootstrap samples were drawn with replacement from the Lichess online database used for the study, each having size the same as the original dataset. The regression was performed on all of the resampled datasets. The bootstrap mean, variances and $95\%$ confidence intervals for the regression coefficients are provided in table 6 and the histograms are presented in the figure 1.
@@
The bootstrap distribution for Chess displays a consistently positive slope across resamples, with a narrow spread around the median. This confirms that increased experience is strongly associated with improved performance, supporting the high skill component in Chess.
5 Analyzing Two Player Rummy
Two Player Rummy is a card game played by 2 players using one or more standard decks of 52 cards (sometimes with jokers). The objective is to form valid combinations of cards called sets (same rank, different suits) or sequences (consecutive cards of the same suit). Each player is dealt a fixed number of cards, and players take turns drawing and discarding to improve their hands. The game incorporates both strategic planning and memory, as players must decide which cards to retain or discard based on observed play. Elements of chance exist through the card draw, but long-term success is strongly influenced by pattern recognition, probability estimation, and timing.
For the analysis of this card game we obtain actual data from online games. We analyze the data for 20,000 players, and the results are in table 7. The results show strong impact of experience in the face of performance. Current experience also has some impact (t-statistic being more than 1).
@@ 5 Analyzing Two Player Rummy
All variables except second-half experience are statistically significant at the 5% level. The first-half win proportion has a particularly strong and significant positive impact, indicating the importance of innate or early-stage skill. First-half experience also shows a positive effect, suggesting that learning through play enhances performance. The model’s modest $R^{2}$ value (0.013) reflects the inherent randomness in card draws and other unobserved factors, but the significance of predictors still supports a meaningful role for skill in Rummy.
5.1 Bootstrap results
The resampling results for Rummy (table 8 and figure 2) show very consistent patterns (very small standard errors and tight support), further supporting the robustness of our skill findings.
@@ 5.1 Bootstrap results
The bootstrap distribution for Rummy shows a consistently positive slope, with most samples indicating a clear upward trend in win rate with experience. The distribution is moderately tight, supporting the statistical robustness of the learning effect observed in the regression. This confirms that skill acquisition plays a substantial role in Rummy, where experienced players tend to perform better over time—despite the inherent randomness in card draws.
6 Analyzing Ludo
Ludo is a multiplayer board game for 2 to 4 players, derived from the Indian game Pachisi. Each player has four tokens that must travel from a starting area to the center of the board by moving around a fixed path. Movement is determined by rolling a six-sided die, and players can capture opponents’ tokens to send them back to the start. While the game is heavily influenced by chance due to dice rolls, strategic choices—such as which token to move and when to block or chase opponents—also play a role. However, the high randomness-to-decision ratio makes it less skill-dominant compared to other games.
@@ 6.1 Ludo Experimental Data
For Ludo, we have used prelaunch experimental data with a small number of players (see Table 9. The regression results are presented in table 10. Here previous performance plays an important role, in addition we also see the number of games in the second tranche being a serious factor with a negative sign. This is probably due to fatigue. The positive contribution of previous experience is also mildly important.
Player ID 1st Half Win % ($w_{1}$) 2nd Half Win % ($w_{2}$) $\phi^{-1}(w_{1})$ $\phi^{-1}(w_{2})$ 1st Half Games ($e_{1}$) 2nd Half Games ($e_{2}$)
@@ 6.2 Ludo Online Data
This study uses data obtained from a new variant of the Ludo game as an integrated offering on the My11Circle (MEC, https://www.my11circle.com/), a fantasy gaming platform. The primary objective of launching the Ludo game on the MEC platform has been to collect the real user game play data. We use this data to conduct our Study. The dataset consists of a total of 265062 players, who have played a total of 3033645 games among themselves during $5^{th}$ June to $23^{rd}$ July. The subsetted dataset used for the actual analysis includes players who have played at least 10 games in both months, consists of a total of 16197 players, who have played a total of 2018935 games, and 1191656 games among themselves during $5^{th}$ June to $23^{rd}$ July.
@@
The regression results in table 11 firmly show that the role of previous experience (learning) and previous performance (innate skill) is quite prominent in determining current performance of a player. This is a strong indicator of the importance of skill in this Ludo game.
6.3 Bootstrap Results
To understand the robustness of our results, we now perform a 10000 fold bootstrap analysis of the regressions results. The consistency of the results through the bootstrap analysis, as evidenced in the bounds of the coefficients in table 12 and tightness of the histograms in figure 3, further strengthen our conclusions.
@@ 6.3 Bootstrap Results
In the case of Ludo, the bootstrap slope distribution is centered around a slightly positive mean but with a relatively wide spread. This indicates that while some users may improve with experience, the association between experience and win rate is weak and highly variable. These findings are in line with the regression results and reflect the game’s high dependence on dice-based randomness, limiting the influence of skill over repeated play.
7 Analyzing Teen Patti
Teen Patti is a gambling-style card game popular in South Asia, often compared to three-card poker. Typically played by 3 to 6 players, each participant is dealt three cards face down. Players bet in rounds, deciding whether to fold, call, or raise based on their hand strength or strategic bluffing. The game has hidden information, high variance, and limited opportunities for strategic depth within a single round. While repeated play may reward probabilistic thinking and bluffing skill, short-term outcomes are heavily driven by chance, especially due to the lack of visibility of other players’ cards.
We consider 5 player games only, as these are the most common. In both the versions, previous performance turns out to be an important factor which probably is also related to innate skills. The learning feature is supported by data as previous experience has a negative impact on current performance. We analyze the two most popular variations of this game, No Limit and Regular.
@@ 7 Analyzing Teen Patti
Table 13) presents the regression summary for the "No Limit" version of Teen Patti, based on a large dataset of over 53,000 users. The model indicates that the first-half win proportion has a strong and highly significant positive effect on current performance, suggesting that initial success—potentially reflective of innate skill or favorable card sequences—plays an important role. Interestingly, the coefficients for first-half experience and second-half experience are significant but of opposite signs: the negative coefficient for first-half experience may suggest overconfidence or fatigue effects, while the positive second-half experience reflects adaptive learning during the session. The model’s $R^{2}=0.116$ shows modest explanatory power, consistent with Teen Patti’s stochastic nature.
Table 14 summarizes the regression results for the "Regular" version of Teen Patti, involving nearly 39,000 users. Here, the first-half win proportion remains a statistically significant positive predictor, reinforcing the importance of initial performance. However, experience-based variables (both first and second half) are statistically insignificant, suggesting little evidence of strategic learning or sustained improvement. This aligns with the perception that Teen Patti outcomes are primarily governed by chance, especially when gameplay is not iterative or memory-based. The low $R^{2}=0.038$ further supports this interpretation, highlighting the difficulty in predicting success from past behavior in this version of the game.
7.1 Bootstrap Results
First we consider the No Limit case. Bootstrap results (see range of estimates in table 15 and histograms in figure 4) confirm the robustness of the regression analysis, indicating that the regression coefficients are likely within the estimated range. The bootstrap estimates exhibit minimal bias, further supporting the reliability of the findings.
@@ 7.1 Bootstrap Results
Table 15: Bootstrap regression summary for 10000 bootstrap estimates (Teen Patti- No Limit)
Figure 4: Histogram of Bootstrap Estimates of Regression Coefficients (Teen Patti-No Limit)
The bootstrap distribution of slope estimates for Teen Patti (No Limit) exhibits a moderate positive skew, indicating some evidence of performance improvement with experience. However, the spread of the distribution is relatively wide, reflecting variability in learning patterns across users. This suggests that while experience can enhance performance in some cases, outcomes remain partially influenced by chance. The results are consistent with the regression findings, highlighting a modest skill component in the No Limit version of Teen Patti.
Next we analyze the Regular one.
@@
Table 16: Bootstrap regression summary for 10000 bootstrap estimates (Teen Patti-Regular)
Figure 5: Histogram of Bootstrap Estimates of Regression Coefficients (Teen Patti-Regular)
The bootstrap slopes for Teen Patti (Regular) (Table16, Figure 5) are centered around zero with a wide spread, indicating weak or no consistent relationship between experience and performance. This supports the interpretation that outcomes in this version are largely driven by chance.
8 Methodology to Generate the Final Skill Score
@@ 8.1 Qualitative Considerations
In this report, our main objective is to draw comparative statements on the effect of luck versus skill on the popular card and board games considered (Tables 4 to 14). The initial discussion revolves around the importance of ”innate skill" (performance in general) and ”learning" (experience: learning by doing). Looking at the comparative results in Table 17 we can assess this for the games under consideration. We certainly look at Chess as the benchmark case. Here rating proxies for skill acquired through experience. The regression results are very clear about the role of previous performance and rating.
Among the other three games, first consider Teen Patti, as this is popularly considered to be the “least skilled” game. Here we see a significant role of “innate skill" but no learning effect. This skill may be attitudinal (bluffing) or cognitive (recognising the behavioural markers of the opponent). But, there is no strategic learning.
Moving on to Rummy the results are very different. Again the relevant ”innate skill" here is the choice of strategy in selection and retention of cards, and there is a significant learning effect. In fact this is true even for current experience (albeit at a lower level of significance). Contrasted with this, if we consider Ludo, we again see both skill as well as learning effect but the effect is now slightly weaker. In fact, the experience effect in Ludo is somewhat mixed with a positive and negative sign.
@@
We see that performance in all the games under consideration rely on innate skill (player heterogeneity), which is one indicator of skill base. Interim findings also suggest that an one shot game like Teen Patti is less amenable to learning compared to the repeated interaction games, e.g. Rummy and the two board games considered. Among the board games, Ludo illustrates a slightly weaker effect of learning compared to Rummy based on the importance of previous experience in terms of the t-stat. See tables 7 and 10, although the difference is very minor. Finally, of course, Chess is the most clear cut skill based game. Based on our limited analysis, we can tentatively rank the four games in order of increasing skill and learning dependence as
@@
However, as we have noted earlier, Ludo and Rummy show very close quantitative scores in terms of our analysis.
@@ 8.2 The Skill Score
We now attempt a scoring exercise on a $[0,1]$ scale for “Skill”. This is tentative as it requires several subjective choices.
@@
$x_{i}=\left\{\begin{array}[]{lll}0&if&t_{i}<0\\ t_{i}/4&if&0\leq t_{i}\leq 2\\ (t_{i}+1)/6&if&2<t_{i}\leq 5\\ 1&if&5<t_{i}\end{array},i=1,2,3.\right.$ (3)
@@
Finally, to define a composite skill score, we propose assigning a weight of 2:1 to the effect of past experience (learning) relative to innate skill. This choice reflects our belief that a player’s ability to improve with repeated play is a more compelling indicator of skill than a high initial win rate alone, which could be due to chance or prior unseen experience. Emphasizing learning aligns with the broader notion of skill as something that develops through mastery, adaptation, and strategic refinement. Moreover, in legal and policy contexts, games are more likely to be classified as skill-based when success can be consistently linked to effort and experience rather than initial performance. In turn, innate skill in the form of past performance should also receive twice as much weight as current experience, as current experience may be confounding learning and fatigue. Thus, we define the skill score as
$s=\frac{2x_{1}+4x_{2}+x_{3}}{7}\in[0,1].$ (4)
While we adopt this weighting in this study, the framework is flexible and can accommodate alternative preferences depending on the context e.g., game design analysis, player behavior modeling, or regulatory thresholds. We hope that we can objectify this choice with a more comprehensive analysis in future work.
Below we illustrate with the games considered here in table 18. The Scores obtained almost confirms our ranking posited above, the only exception being Ludo scores slightly higher than Rummy now. As these two numbers are again very close to each other, we think that the results may be sensitive to the choice of cut-offs and weightages in the score formula.
@@ 8.3 Resampling Results on Skill Score
The results are not only dependent on the choice of the score function. Sampling fluctuation inherent in the data may also influence our scores and hence ranking. So we also carry out a bootstrap analysis with our regression model. We will get a range of t-stat based on the range of regression coefficients from bootstrap analysis. This will give us a range of $t_{i}$ values and hence a range of $x_{i}$ values. Thus, instead of a point estimate for the score function $s$ for each game, it will have a range of values. We can then compare distribution of the score values to arrive at a comprehensive ranking template. The details are available in the following table.
@@
For each of the games we use 10000 bootstrap samples to generate the t-statistics. Hence, we get a distribution of scores for each of the combinations in table 19. These are presented in figures 6 for Chess, 7 for Rummy, 8 for Ludo and finally 9 for Teen Patti at the end of this paper.
@@
To simplify our discussion, we produce the median and an 80% confidence interval for all the score distributions in table 21 where the degeneracy of Chess score is clear. It is also clear that Teen Patti is less skilled than either Rummy or Ludo (the confidence intervals are practically separated) who are in turn less skilled than Chess. The comparison between Rummy and Ludo is unclear, as already guessed at earlier. The distributions overlap with Ludo having a little more spread and the centre of the distributions (median) also switching order between Rummy and Ludo for alternative parametric configurations. This clearly indicates that these two games are very similar in total skill content as indicated earlier. Of course, the kind of skill required may be quite diverse.
@@
9 Concluding Remarks
It is important to note that some of our regression models yield relatively low $R^{2}$ values. However, this should not be viewed as undermining the validity of the models or the significance of the predictors. In large-scale observational datasets, especially those capturing human behavior such as game performance outcomes are often influenced by numerous unobserved or idiosyncratic factors (e.g., mood, device used, time of day), which contribute to the residual variance. A low $R^{2}$ in such contexts does not imply that the identified predictors are uninformative; rather, it reflects the inherent noise in the data and the partial nature of the model’s explanatory scope. In fact, when sample sizes are large and predictors are theoretically grounded, the significance of regression coefficients (e.g., p-values, t-statistics, and bootstrap confidence intervals) often offers a more meaningful interpretation than the overall variance explained. This is especially common in behavioral and social science data, where $R^{2}$ values as low as 0.01–0.05 are considered acceptable if effect sizes are statistically and practically significant [14],[15]. Our use of resampling-based confidence intervals further reinforces the robustness of our coefficient estimates, even when $R^{2}$ is modest.
The conclusions from our analysis are quite nuanced with Chess being the most skilled (one extreme) and Teen Patti on the other end (but not an extreme score, around 0.45). While Rummy and Ludo are in between, the ranking between them is ambiguous. Depending on cutoffs and weight-age schemes, the score distribution is always overlapping for these two with a wider range for Ludo than Rummy (possibly due to the larger sample of Rummy data).
While the final skill score relies on subjective choices of cut-offs and weights, we provide multiple formulations to demonstrate robustness. This flexibility is intentional, as different stakeholders (e.g., legal, gaming design, regulatory) may prioritize components like learning rate or innate performance differently. Thus, the framework is not meant to offer a single "correct" number, but rather a statistically grounded range to inform debate. Thus, as mentioned in the Introduction, we allow for a difference of opinion regarding the skill comparison of Ludo and Rummy. Their skill profiles are unterminated by the other and hence their ranking will change with a change in scaling and weight choice. We feel that this provides a simple but flexible data-driven comparison tool for skill games in general.
While our current framework is built on win statistics, it can be extended to accommodate other definitions of successful gameplay. For example, in cooperative or resource-based games, success may involve minimizing loss, optimizing resources, or sustaining high performance across rounds. Incorporating additional performance metrics would allow for a richer, game-specific evaluation of skill. This is a promising direction for future work.